# Coherence-Assisted Entanglement Activation During Open Evolution

### v2

@@ 3 Bounds on the system-environment entanglement generation
Let us look closely at Lemma 3 and give some interpretations. The second term $H(i|m)$ on the right hand side is the uncertainty about the system’s state measured by the environment after our best measurement on environment. Non zero $H(i|m)$ means that we cannot conclude the system’s value in the pointer basis even if we could use any POVM on the environment. Reversely, if $H(i|m)$ becomes zero, the information about the system is transferred fully to environment and entanglement must be there unless system’s coherence is zero, which means the system is fully classical state. After all, when the system has more coherence in pointer bases, more entanglement is required to know the value of system certainly. One can also re-arrange the inequality (13) and obtain the lower bound on the uncertainty about the system’s state in terms of initial coherences and the generated entanglement:
$\max_{POVM:\{M_{m}\}}H(i|m)\geq C_{r}(\rho_{S})-E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger}).$ (22)
Similar interpretation can be applied to the qubit version (1). If $p(1-p)=0$ one knows the system’s state with certainty. Notice that $C_{r}(\rho_{S})=0$ in this case. Also if $B(\rho_{0}(t),\rho_{1}(t))=0$, it implies that $\rho_{0}(t)$, $\rho_{1}(t)$ are perfectly distinguishable, which means maximum information has been transferred to the environment.
@@
Finally, we will give an upper bound and another lower bound which we mentioned in the proof of lemma 3 on the generated entanglement. From a previous research [17], an upper bound on entanglement generated by incoherent operations with initially coherent system $\rho_{S}$ and an incoherent pure ancilla $|0\rangle\langle 0|_{E}$ is given by the initial coherence in system, for any measure of entanglement $E_{D}$ and coherence $C_{D}$ corresponding to a contractive distance $D$:
@@
However, since the corresponding unitary evolution $U$ for given Hamiltonian of (3) is not incoherent in general regardless of incoherent basis for environment and we assume the environmental state arbitrary, the above upper bound is not valid in our scenario. Also, considering our intuitive picture of system-environment interaction where coherence is converted into entanglement, it is interesting to connect the generated entanglement not only to the initial coherence $C_{r}(\rho_{S})$ as we did in lemma 3 for lower bound, but to the decoherence, the difference of coherence between the initial and final states in system as we demonstrate below.
@@
Let $D_{S}$ be the decoherence in the system defined as $D_{S}=C_{r}(\rho_{S})-C_{r}(\sigma_{S})$. Then the relative entropy of entanglement generated by pure dephasing evolution $U$ is bounded by below and above as:
$D_{S}-S(\rho_{E})\leq E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})\leq D_{S}+H(i).$ (24)
@@
Let us prove the left inequality first. From lemma 1, $E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})\geq S(\sigma_{S})-S(\sigma_{SE})$. The invariancy of entropy over unitary evolution and the fact $S(\Delta(\sigma_{S}))=S(\Delta(\rho_{S}))$ lead us to the followings.
$E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})\geq S(\sigma_{S})-S(\sigma_{SE})$
$=$ $S(\sigma_{S})-S(\rho_{S})-S(\rho_{E})$
$=$ $S(\sigma_{S})-S(\Delta(\rho_{S}))+S(\Delta(\rho_{S}))-S(\rho_{S})-S(\rho_{E})$
$=$ $C_{r}(\rho_{S})-C_{r}(\sigma_{S})-S(\rho_{E})$ (25)
Moving to the right hand side, from the definition of the relative entropy of entanglement: $E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})\leq S(\sigma_{SE}||\sigma_{S}\otimes\sigma_{E})=I(S:E)=S(\sigma_{S})+S(\sigma_{E})-S(\sigma_{SE})=S(\sigma_{S})+S(\sigma_{E})-S(\rho_{S})-S(\rho_{E})$, since by definition $\sigma_{SE}=U\rho_{S}\otimes\rho_{E}U^{\dagger}$ so that $S(\sigma_{SE})=S(\rho_{S}\otimes\rho_{E})=S(\rho_{S})+S(\rho_{E})$. We also observe that $S(\sigma_{E})=S(\sum_{i}p_{i}\rho_{i})\leq S(\rho_{E})+H(i)$, which follows from the well known property of the von Neumann entropy: $S(\sum_{i}p_{i}\rho_{i})\leq\sum p_{i}S(\rho_{i})+H(i)$. Substituting this into the above inequality, we prove the statement. ∎
@@
$E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})-H(i)\leq D_{S}.$ (26)
4 Case study: The spin-boson model
@@
$H$ $=\sum_{\boldsymbol{k}}\omega_{\boldsymbol{k}}a^{\dagger}_{\boldsymbol{k}}a_{\boldsymbol{k}}+\sigma_{z}\otimes\sum_{\boldsymbol{k}}\left(g_{\boldsymbol{k}}a^{\dagger}_{\boldsymbol{k}}+g^{*}_{\boldsymbol{k}}a_{\boldsymbol{k}}\right)$ (27)
where for the purpose of this example, we have suppressed the self-Hamiltonian of the central system, or assumed it to be commuting with the interaction, as otherwise the evolution would not be of the pure dephasing form (4). Here, $a^{\dagger}_{\boldsymbol{k}}$, $a_{\boldsymbol{k}}$ are the annihilation and creation operators of the mode $\boldsymbol{k}$ of the environment, $\sigma_{z}$ is the $z$ axis Pauli matrix acting in the space of the central system and $g_{\boldsymbol{k}}$ are the coupling constants. The pointer basis here is the eigenbasis of $\sigma_{z}$, which follows immediately from the interaction term in (27).
@@
$\frac{1}{2}\ln B_{vac}(t)=$ $-\wp(s-1)\left[1-\frac{\cos[(s-1)\arctan(\Lambda t)]}{(1+\Lambda^{2}t^{2})^{\frac{s-1}{2}}}\right],$ (32)
$\frac{1}{4}\ln B_{th}(t)=$ $\left(-\frac{T}{2\Lambda}\right)^{s-1}\bigg{\{}\Psi^{(s-2)}\left(1+\frac{T}{2\Lambda}\right)-\Psi^{(s-2)}\left(\frac{1}{2}+\frac{T}{2\Lambda}\right)$
@@
where c.c. means complex conjugate of the last two terms in the parenthesis, $\wp(s)$ is the Euler gamma function, and $\Psi^{m}(z)$ is the polygamma function defined as:
$\Psi^{m}(z)=\sum_{k=0}^{\infty}\frac{(-1)^{m+1}m!}{(z+k)^{m+1}}.$ (34)
Figure 1: Lower bound (1) for the spin-boson model as a function of time for a different mixedness parameter $\alpha$ (cf. (28)) and different Ohmicity: a) $s=2$ corresponding to Markovian evolution; b) $s=3$ corresponding to non-Markovian evolution. The ratio of the thermal to the cutoff energy is set to $T/\Lambda=1$ and the time is measured in the cutoff units. The insert shows the peak and thus the non-monotonic behavior of the plots (with different vertical scale for each plot). On all plots the parts, where the bound becomes negative were replaced by the trivial bound equal to zero.
[…]
In Fig. 1, we keep fixed the ratio of the thermal to the cutoff energy, $T/\Lambda=1$, and plot (1) for different mixedness $\alpha$ and Ohmicity $s$ parameters. As expected from the (1), less mixed, i.e. more coherent, states with lower $\alpha$ lead to much more entanglement generation. More interestingly, there is a slight quantitative change of behavior between $s=2$ and $s=3$, corresponding to the transition between the Markovian and non-Markovian evolution respectively. While for $s=2$ the behavior is monotonic in time, for $s=3$ there is a broad peak around $\Lambda t=2$, regardless the value of $\alpha$ (depicted on the insert). This peak is more visible in Fig. 2, where we fix the mixedness $\alpha=1/2$ and vary the energy ratio $T/\Lambda$. The non-monotonic behavior of the entanglement bound for $s=3$ bears certain similarity to non-Markovianity measures, where a non-monotonic behavior of some marker function certifies information back-flow and hence the non-Markovianity of the evolution. Especially close seems the measure from [36], where the non-Markovianity is detected via non-monotonicity of the decoherence rate, which in turn derives from system-environment correlations.
@@
5 Conclusions
We studied generation of entanglement between the system of interest and its environment for a broad class of so called pure dephasing evolutions, also known as measurement-type interactions. In this part of the work, we extended previous results  [4, 5] obtained for pure initial states of the system, to the case of mixed states. We combined methods and results from coherence theory, state discrimination, and open systems evolution to obtain a series of bounds on the amount of the generated entanglement as measured by the relative entropy. Firstly, and most importantly, we obtain lower bounds, which certify that at least the given amount of entanglement was created during the evolution out of a given amount of the initial coherence. The lower bound character of this result is a price to pay for the generalization to arbitrary, mixed states of the system. For qubit as the central system, where we obtained a closed expression, connecting amount of entanglement to the state distinguishability of the environmental states. This makes an interesting link to certain studies of quantum Darwinism [9]. For an arbitrary dimension, our bound still contains a certain optimization. As an example, we calculated analytically and studied numerically the lower bound for the popular spin-boson system, rediscovering the well-known traces of the non-Markovianity but in the entanglement dynamics. Obviously, more systems can be analyzed, especially that for some of them the relevant state distinguishability has already been quantified [37, 38, 39, 33].
Next, we derived both lower and upper bounds for the entanglement in terms of the difference of coherence in the system before and after the evolution, which is a measure of the decoherence efficiency. This proves in an elegant and general way that stronger entanglement leads to more decoherence and vice versa, more decoherence is accompanied with more entanglement generation. These bounds also can be seen as an extension of the result obtained in [17].
Summarizing, our results allow to better characterize and understand the correlations that pure dephasing interaction can produce under general initial conditions. Because of the ubiquitous character of the pure dephasing interaction in open systems [2, 3], our results are of a fairly universal character across the field of open systems and decoherence theory. Moreover, since quantum measurement theory is largely based on the same type of interaction, we characterize how much entanglement is generated during a measurement process, thus linking to foundations of quantum mechanics. On the other hand, a possible practical importance of our studies may lie in the development of non-standard decoherence-mitigation techniques, which rely on the system-environment correlations (see e.g. [40] for an example of such a technique).

### v5

@@ III Bounds on the system-environment entanglement generation
Let us look closely at Lemma 3 and give some interpretations. The second term $H(\mathcal{I}|\mathcal{M})$ on the right hand side is the uncertainty about the system’s state measured by the environment after our best measurement on environment. Non zero $H(\mathcal{I}|\mathcal{M})$ means that we cannot conclude the system’s value in the pointer basis even if we could use any POVM on the environment. Conversely, if $H(\mathcal{I}|\mathcal{M})$ becomes zero, the information about the system is transferred fully to environment and entanglement must be there unless system’s coherence is zero, which means the system is fully classical state. After all, when the system has more coherence in pointer bases, more entanglement is required to know the value of system certainly. One can also re-arrange the inequality (14) and obtain the lower bound on the uncertainty about the system’s state in terms of initial coherences and the generated entanglement:
$\max_{POVM:\{M_{m}\}}H(\mathcal{I}|\mathcal{M})\geq C_{r}(\rho_{S})-E_{r}(\sigma_{SE}).$ (22)
Similar interpretation can be applied to the qubit version (21). If p(1p)=0p(1-p)=0 one knows the system’s state with certainty. Notice that $C_{r}(\rho_{S})=0$ in this case. Also if $B(\rho_{0}(t),\rho_{1}(t))=0$, it implies that $\rho_{0}(t)$, $\rho_{1}(t)$ are perfectly distinguishable, which means maximum information has been transferred to the environment.
@@
Finally, we will give an upper bound on the generated entanglement. From a previous research [17], an upper bound on entanglement generated by incoherent operations with the initially coherent system $\rho_{S}$ and an incoherent pure ancilla $|0\rangle\langle 0|_{E}$ is given by the initial coherence in system, for any measure of entanglement $E_{D}$ and coherence $C_{D}$ corresponding to a contractive divergence $D$:
@@
However, since the corresponding unitary evolution $U$ for the given Hamiltonian of (3) is not incoherent in general regardless of incoherent basis for environment and we assume the environmental state arbitrary, the above upper bound is not valid in our scenario. Also, considering our intuitive picture of system-environment interaction where coherence is converted into entanglement, it is interesting to connect the generated entanglement not only to the initial coherence $C_{r}(\rho_{S})$ as we did in lemma 3 for lower bound, but to the decoherence $C_{r}(\rho_{S})-C_{r}(\sigma_{S})$, the difference of coherence between the initial and final states in system as we demonstrate below.
@@
The relative entropy of entanglement generated by pure dephasing evolution $U$ is bounded from above as:
$E_{r}(\sigma_{SE})\leq C_{r}(\rho_{S})-C_{r}(\sigma_{S})+H(\mathcal{I}).$ (24)
@@
From the definition of the relative entropy of entanglement, $E_{r}(\sigma_{SE})\leq S(\sigma_{SE}||\sigma_{S}\otimes\sigma_{E})$. We observe that $\log\sigma_{S}\otimes\sigma_{E}=\log\sigma_{S}\otimes\mathbbm{I}+\mathbbm{I}\otimes\log\sigma_{E}$ and $\textrm{Tr}[\sigma_{SE}\log\sigma_{S}\otimes\mathbbm{I}]=\textrm{Tr}_{S}[\textrm{Tr}_{E}[\sigma_{SE}]\log\sigma_{S}]=\textrm{Tr}_{S}[\sigma_{S}\log\sigma_{S}]=-S(\sigma_{S}).$ Then we have
$E_{r}(\sigma_{SE})\leq$ $\ S(\sigma_{SE}||\sigma_{S}\otimes\sigma_{E})$
$=$ $-S(\sigma_{SE})-\textrm{Tr}[\sigma_{SE}\log\sigma_{S}\otimes\sigma_{E}]$
$=$ $-S(\rho_{S})-S(\rho_{E})-S(\sigma_{S})+S(\sigma_{E})$ (25)
We also observe that $S(\sigma_{E})\leq S(\rho_{E})+H(\mathcal{I})$, which follows from the well known property of the von Neumann entropy: $S(\sum_{i}p_{i}\rho_{i})\leq\sum p_{i}S(\rho_{i})+H(\mathcal{I})$. Substituting this into the above inequality, we prove the statement. ∎
@@
$E_{r}(\sigma_{SE})-H(\mathcal{I})\leq C_{r}(\rho_{S})-C_{r}(\sigma_{S}).$ (26)
IV Case study: The spin-boson model
@@
$H$ $=\sum_{\bm{k}}\omega_{\bm{k}}a^{\dagger}_{\bm{k}}a_{\bm{k}}+\sigma_{z}\otimes\sum_{\bm{k}}\left(g_{\bm{k}}a^{\dagger}_{\bm{k}}+g^{*}_{\bm{k}}a_{\bm{k}}\right)$ (27)
where for the purpose of this example, we have suppressed the self-Hamiltonian of the central system, or assumed it commutes with the interaction, as otherwise the evolution would not be of the pure dephasing form (4). Here, $a^{\dagger}_{\bm{k}}$, $a_{\bm{k}}$ are the annihilation and creation operators of the mode $\bm{k}$ of the environment, $\sigma_{z}$ is the $z$ axis Pauli matrix acting in the space of the central system and $g_{\bm{k}}$ are the coupling constants. The pointer basis here is the eigenbasis of $\sigma_{z}$, which follows immediately from the interaction term in (27).
@@
$\frac{1}{2}\ln B_{vac}(t)=$ $-\Gamma(s-1)\left[1-\frac{\cos[(s-1)\arctan(\Lambda t)]}{(1+\Lambda^{2}t^{2})^{\frac{s-1}{2}}}\right],$ (32)
$\frac{1}{4}\ln B_{th}(t)=$ $\left(-\frac{T}{2\Lambda}\right)^{s-1}\bigg\{\Psi^{s-2}\left(1+\frac{T}{2\Lambda}\right)-\Psi^{s-2}\left(\frac{1}{2}+\frac{T}{2\Lambda}\right)+\left[\frac{1}{2}\Psi^{s-2}\left(\frac{1}{2}+\frac{T}{2\Lambda}+i\frac{Tt}{2}\right)-\frac{1}{2}\Psi^{s-2}\left(1+\frac{T}{2\Lambda}+i\frac{Tt}{2}\right)+\textrm{c.c.}\right]\bigg\},$ (33)
where c.c. means complex conjugate of the two terms in the square parenthesis, $\Gamma(s)$ is the Euler gamma function, and $\Psi^{m}(z)$ is the polygamma function defined as:
$\Psi^{m}(z)=\sum_{k=0}^{\infty}\frac{(-1)^{m+1}m!}{(z+k)^{m+1}}.$ (34)
@@
Figure 2: The same as in Fig. 1 but for a fixed mixedness parameter $\alpha=1/2$ and different ratios $T/\Lambda$. a) Ohmicity $s=2$; b) Ohmicity $s=3$.
The resulting behavior of the lower bound is presented in Figs. 1, 2. The regions in the plots where the right-hand side of (21) becomes negative, i.e. the bound trivializes since by definition $E_{r}\geq 0$ always, were substituted by zero to: i) avoid a suggestion that $E_{r}$ might be negative; ii) emphasize what might be called "a sudden revival of the bound". Of course, since this is a lower bound, the fact that in some situations the bound trivializes does not mean there cannot be any entanglement and thus the discontinuity for the initial times need not mean a sudden birth of entanglement. But on the other hand, it is intuitively appealing that since we initiate the whole model in a product state, it will take some time for the central system and the environment to get entangled.
In Fig. 1, we keep fixed the ratio of the thermal to the cutoff energy, $T/\Lambda=1$, and plot (21) for different mixedness $\alpha$ and Ohmicity $s$ parameters. As expected from the (21), less mixed, i.e. more coherent, states with lower $\alpha$ lead to much more entanglement generation. More interestingly, there is a slight quantitative change of behavior between $s=2$ and $s=3$, corresponding to the transition between the Markovian and non-Markovian evolution respectively. While for $s=2$ the behavior is monotonic in time, for $s=3$ there is a broad peak around $\Lambda t=2$, regardless of the value of $\alpha$ (depicted on the inset). This peak is more visible in Fig. 2, where we fix the mixedness $\alpha=1/2$ and vary the energy ratio $T/\Lambda$. The non-monotonic behavior of the entanglement bound for $s=3$ bears certain similarity to non-Markovianity measures, where a non-monotonic behavior of some marker function certifies information back-flow and hence the non-Markovianity of the evolution. Especially close seems the measure from [38], where the non-Markovianity is detected via non-monotonicity of the decoherence rate, which in turn derives from system-environment correlations.
@@
V Conclusions
We studied generation of entanglement between the system of interest and its environment for a broad class of so-called pure dephasing evolutions, also known as measurement-type interactions. In this part of the work, we extended previous results  [4, 5] obtained for pure initial states of the system, to the case of mixed states. We combined methods and results from coherence theory, state discrimination, and open systems evolution to obtain a series of bounds on the amount of the generated entanglement as measured by the relative entropy. Firstly, and most importantly, we obtained a lower bound, which certifies that at least the given amount of entanglement was created during the evolution out of a given amount of the initial coherence. The lower bound character of this result is a price to pay for the generalization to arbitrary, mixed states of the system. For a qubit as the central system, we obtained a closed expression that connects the amount of entanglement to the state distinguishability of the environmental states. This makes an interesting link to certain studies of quantum Darwinism [9]. For an arbitrary dimension, our bound still contains a certain optimization. As an example, we calculated analytically and numerically the lower bound for the popular spin-boson system, rediscovering well-known traces of the non-Markovianity in the entanglement dynamics. Clearly, more systems can be analyzed, especially since for some of them the relevant state distinguishability has already been quantified [39, 40, 41, 35].
Next, we derived an upper bound for the entanglement in terms of the difference of coherence in the system before and after the evolution, which is a measure of the decoherence efficiency. This proves in an elegant and general way that stronger entanglement leads to more decoherence and vice versa, more decoherence is accompanied by more entanglement generation. This bound can also be seen as an extension of the result obtained in [17].
Summarizing, our results allow us to better characterize and understand the correlations that pure dephasing interaction can produce under general initial conditions. Because of the ubiquitous character of the pure dephasing interaction in open systems [2, 3], our results are fairly universal across the field of open systems and decoherence theory. Moreover, since quantum measurement theory is largely based on the same type of interaction, we characterize how much entanglement is generated during a measurement process, thus linking to the foundations of quantum mechanics. On the other hand, a possible practical importance of our studies may lie in the development of non-standard decoherence-mitigation techniques, which rely on the system-environment correlations (see e.g. [42] for an example of such a technique).